Calculate the mass of a solid iron sphere that has a diameter of 3.0 cm.

1
Expert's answer
2021-12-15T11:55:25-0500

"m=\rho\cdot V=\rho\cdot\frac{4}{3}\cdot\pi\cdot r^3=7874\cdot\frac{4}{3}\cdot3.14\cdot 0.03^3="


"=0.89\ (kg)=890\ (g)" . Answer